OpenStudy (anonymous):

Jarod borrowed $500 from his parents. His parents are charging a simple interest rate of 3%. How much interest will Jarod pay if he pays his parents back in 6 months? Will give medal :)

OpenStudy (anonymous):

How about this answer? A = $ 507.50 Equation: A = P(1 + rt) Calculation: First, converting R percent to r a decimal r = R/100 = 3%/100 = 0.03 per year, putting time into years for simplicity, 6 months ÷ 12 months/year = 0.5 years, then, solving our equation A = 500(1 + (0.03 × 0.5)) = 507.5 A = $ 507.50 The total amount accrued, principal plus interest, from simple interest on a principal of $ 500.00 at a rate of 3% per year for 0.5 years (6 months) is $ 507.50. Found it here - http://www.calculatorsoup.com/calculators/financial/simple-interest-plus-principal-calculator.php